# FreqKV: Frequency Domain Key-Value Compression for Efficient Context Window Extension

### v2

@@ Abstract
Frequency-domain compression has proven effective in reducing redundancies for spatial signals. In this work, we propose FreqKV, a novel frequency domain key-value (KV) compression technique that enables efficient context window extension for decoder-only large language models (LLMs). Our approach is motivated by a key observation that, in the frequency domain, the energy distribution of the KV cache is predominantly concentrated in low-frequency components. By discarding high-frequency components, we achieve efficient compression of the KV cache with minimal information loss. FreqKV iteratively compresses the increasing KV cache to a fixed size in the frequency domain, allowing models to process lengthy contexts efficiently. Introducing no additional parameters or architectural modifications, FreqKV is applicable to both fine-tuning and inference. With minimal fine-tuning, LLMs can learn to leverage the limited cache that is compressed in the frequency domain and extend the context window. Experiments on a range of long context language modeling and understanding tasks demonstrate the efficiency and effectiveness of the proposed method.
@@ 1 Introduction
Large language models (LLMs) typically have a limited size of context window, which is pre-defined during the pre-training process. However, LLMs struggle to maintain their performance when processing sequences that exceed the preset context size. Additionally, the computational cost of the self-attention mechanism (Vaswani et al.,, 2023) grows quadratically with the context length, meaning that doubling the context window results in a fourfold increase in the computational cost of attention modules.
Recent studies propose to fine-tune LLMs to longer contexts to extend the context window. LongLoRA (Chen et al.,, 2024) trains LLMs using shifted sparse attention. Despite training efficiency, their sparse attention fails to be applied during inference, and they still require the original attention on the full sequence. Concurrently, LoCoCo (Cai et al., 024a, ) and Activation Beacon (Zhang et al., 2024a, ) introduce additional modules to compress KV states and incorporate the fine-tuned compressing pattern into the decoding procedure of LLMs. However, the increasing memory requirement for additional parameters is inevitable.
For efficiency, existing efforts aim to compress the key-value (KV) cache for long contexts during inference. They evict (Xiao et al.,, 2024; Li et al.,, 2024; Jo et al.,, 2025) or merge (Zhang et al., 2024b, ; Wan et al.,, 2024) KV states of less important tokens following certain rules. They use attention scores to measure the importance and approximate the original full attention. However, while these methods provide an approximation of full computation on existing tokens, the information associated with the evicted tokens will be lost. They can not fully prevent performance degradation when decoding future tokens.
In the field of computer vision, studies have shown that low-frequency channels are more important for convolutional neural networks (CNNs) (Xu et al.,, 2020). Moreover, Fourier Transformer (He et al.,, 2023) discards the high-frequency parts of the contexts and downsample the hidden states in the encoders. Inspired by these works, we seek to compress KV states in the frequency domain without the need for additional compression modules in decoder-only LLMs. We transform key states and value states in LLaMA-2-7b (Touvron et al.,, 2023) from the time domain to the frequency domain for power spectrum analysis. As shown in Figure 1, the energy distribution is increasingly concentrated on low-frequency components as the computation process progresses along the decoder layers. It suggests that high-frequency components, which occupy less energy and contribute less to the overall information, can be discarded without much performance loss, thereby enhancing computational efficiency.
In this paper, we introduce FreqKV, an efficient context window extension method that iteratively compresses key-value states in the frequency domain. Compression is triggered only when the KV cache reaches the predefined context window size. During each compression step, KV states are converted to the frequency domain, where low-frequency components are retained before being transformed back to the time domain. Subsequent tokens are appended to the compressed cache until it is filled again. This ensures that the cached KV states that each query token can attend to are limited below the context window size, reducing memory and computational costs. The compressed cache undergoes further compression alongside incoming tokens. This iterative process progressively increases the compression ratio of earlier contexts as the sequence lengthens, while recent tokens remain less compressed, which is beneficial for autoregressive models to predicet next tokens. Without introducing additional compression modules, LLMs could learn to utilize the compressed cache efficiently when extending to longer contexts. FreqKV demonstrates comparable performance to other methods that employ full KV cache or additional compressors in long context language modeling. Furthermore, experiments on downstream tasks indicate that FreqKV surpasses recently studied KV compression methods in long context understanding.
@@ 2 Related Work
Context Extension for LLMs.
Recent advancements in context extension for LLMs have focused on efficiently scaling models to handle longer input sequences without significantly increasing computational costs. LongLoRA (Chen et al.,, 2024) employs shifted sparse attention during the parameter-efficient fine-tuning. However, this sparse attention mechanism is not applicable during inference, necessitating a return to the original full attention post-training. Other techniques, such as LoCoCo (Cai et al., 024a, ), integrate convolutional operations into LLMs for compressing long contexts. They fine-tune the compression modules together with LLMs. Landmark attention (Mohtashami and Jaggi,, 2023) uses landmark tokens to retrieve previous input blocks. Similarly, Activation Beacon (Zhang et al., 2024a, ) introduces a special token to represent the previous context for compression. However, they need a copy of multi-head attention, which amounts to approximately 2B for 7B models. In contrast, our proposed method achieves context extension without introducing any additional parameters.
@@
To extend the context window of LLMs efficiently, researchers attempt to compress the KV cache as more tokens are fed into the model. One common approach is selective token eviction (Xiao et al.,, 2024; Li et al.,, 2024; Cai. et al., 024b, ; Jo et al.,, 2025), where less significant tokens are discarded. Although the eviction strategies ensure that the size of KV cache involved in each decoding step does not exceed the pre-defined context window size, LLMs suffer from the permanent loss of the information associated with evicted tokens. To address this limitation, some researchers introduce cache merging techniques to approximate the original full attention of the existing contexts (Zhang et al., 2024b, ; Wan et al.,, 2024; Wang et al.,, 2024). However, these inference methods often sacrifice performance for efficiency.
@@
Learning in the frequency domain is a well-established technique to compress images and accelerate CNNs (Gueguen et al.,, 2018). It has been observed that CNNs are more sensitive to low-frequency channels than high-frequency channels (Xu et al.,, 2020). These works have inspired efforts to process natural language. FNet (Lee-Thorp et al.,, 2022) improves the efficiency of Transformer encoder architectures by replacing the self-attention layers with the Fourier transform to serve the purpose of mixing tokens. Additionally, Fourier Transformer (He et al.,, 2023) eliminates redundancies in the context through frequency domain processing within encoder architectures.
@@
3 Preliminaries
3.1 Discrete Cosine Transform
$y_{t}=\alpha_{t}\sum_{n=0}^{N-1}x_{n}\cdot\cos\left[\frac{\pi t(2n+1)}{2N}\right],$ $\alpha_{t}=\begin{cases}\sqrt{\frac{1}{N}}&\text{ if }t=0,\\ \sqrt{\frac{2}{N}}&\text{ otherwise }\end{cases}$ (1)
where $t=0,1,\cdots,N-1$. $\alpha_{t}$ is the normalization factor.
The time-domain signal ${\bm{X}}_{0:N-1}$ can be recovered by applying the inverse DCT (IDCT) on the frequency components ${\bm{Y}}_{0:N-1}$:
$x_{n}=\sum_{t=0}^{N-1}\alpha_{t}\cdot y_{t}\cdot\cos\left[\frac{\pi t(2n+1)}{2N}\right].$ (2)
The frequency components are expressed as a combination of the original signals. The values can be computed using the Fast Fourier Transform (FFT) with a complexity of $O(N\text{log}N)$. The amplitudes of frequency components are utilized in the power spectrum analysis to represent the energy or magnitude of components. The components of higher energy in the frequency domain indicate they are more informative (He et al.,, 2023).
3.2 Self-Attention
For the incoming token $x_{N}$, the prefilled $N$ tokens ${\bm{X}}_{0:N-1}$ are utilized as the cache during decoding. Denote the cached KV states for the previous $N$ tokens ${\bm{X}}_{0:N-1}$ as ${\bm{K}}_{0:N-1}$ and ${\bm{V}}_{0:N-1}$. For simplicity, indices corresponding to layers and heads have been omitted.
@@ 4 FreqKV
Figure 1: The average power spectrums of key states and value states in different layers of LLaMA-2-7b. 1000 documents are sampled from CNN/Daily Mail (Hermann et al.,, 2015). We use DCT to transform key states and value states to the frequency domain, and average power spectrums over these samples and hidden dimensions.
4.1 Energy Concentration in the Frequency Domain
We transform key states and value states from the time domain, which is the sequence dimension, to the frequency domain. The average power spectrums in different decoder layers of LLaMA-2-7b are calculated and presented in Figure 1.
The figure shows that the energy of key states and value states is increasingly concentrated in the low-frequency components. The distributions of their power spectrums are similar since they are projected from the same hidden states. Although there is no obvious energy concentration in the frequency domain for the initial embeddings of natural languages in the first layer, the model tends to aggregate energy in the low-frequency components along the decoding procedure. The observation of energy concentration suggests that we could maintain low-frequency components and filter out high-frequency components which could be redundant. Head-wise analysis of the power spectrum is provided in Appendix A.
@@ 4.2 KV Compression in the Frequency Domain
We conduct DCT along the sequence dimension to transfer the KV cache to the frequency domain:
${\bm{Y}}_{0:N-1}^{K}=\text{DCT}(\displaystyle{\bm{K}}_{0:N-1}),\quad\displaystyle{\bm{Y}}_{0:N-1}^{V}=\text{DCT}(\displaystyle{\bm{V}}_{0:N-1}).$ (4)
As observed in Figure 1, since the lower-frequency components are of higher magnitude and carry more information, we will retain them and remove higher-frequency components for compression. Given the retaining ratio $\gamma$, the retaining size is $L=\gamma\cdot N$. $N-L$ high-frequency components are filtered out to reduce redundancy:
$\widetilde{\displaystyle{\bm{Y}}}_{0:L-1}^{K}=\displaystyle{\bm{Y}}_{0:N-1}^{K}[0:L-1],\quad\widetilde{\displaystyle{\bm{Y}}}_{0:L-1}^{V}=\displaystyle{\bm{Y}}_{0:N-1}^{V}[0:L-1].$ (5)
Then we conduct IDCT along the frequency dimension to convert the compressed components back to the time dimension. It should be noted that the signals are normalized by the square root of the component number as shown in the formulas of DCT and IDCT. While DCT works on $N$ time-domain signals, only $L$ frequency-domain signals are processed by IDCT. The compressed states are amplified by $\sqrt{\frac{N}{L}}$ during the tranforms. Therefore, the compressed signals should be rescaled with the factor of $\sqrt{\frac{L}{N}}$ to restore the original amplitude:
$\widetilde{\displaystyle{\bm{K}}}_{0:L-1}^{0:N-1}=\sqrt{\frac{L}{N}}\text{IDCT}(\widetilde{\displaystyle{\bm{Y}}}_{0:L-1}^{K}),\quad\widetilde{\displaystyle{\bm{V}}}_{0:L-1}^{0:N-1}=\sqrt{\frac{L}{N}}\text{IDCT}(\widetilde{\displaystyle{\bm{Y}}}_{0:L-1}^{V}).$ (6)
$\widetilde{\displaystyle{\bm{K}}}_{0:L-1}^{0:N-1}$ and $\widetilde{\displaystyle{\bm{V}}}_{0:L-1}^{0:N-1}$ are the compressed KV cache of size $L$ in the time domain. The superscript “0:$N-1$” means that $\widetilde{\displaystyle{\bm{K}}}_{0:L-1}^{0:N-1}$ and $\widetilde{\displaystyle{\bm{V}}}_{0:L-1}^{0:N-1}$ are the compressed KV of the cached $N$ tokens. The subscript “0:$L-1$” means the retaining size is $L$. The incoming token $x_{N}$ will attend to the compressed KV cache:
@@
Figure 2: The illustration of our FreqKV. The KV cache will be compressed in an iterative manner as the cache reaches the context window size. Sink tokens remain uncompressed throughout the process. The tokens after sink tokens will be compressed in the frequency domain and subsequent tokens will continue to get into the cache. When the cache is filled again, the compressed tokens and incoming tokens will be compressed together. The compression is performed iteratively to extend the context window.
@@ 4.3 Context Extension via Iterative Compression
When extending the context window, the memory requirement of KV cache increases linearly and the computation cost of the original full attention grows quadratically with the length. To limit the size of KV cache for each query token to attend below the context window size, we will compress KV cache in the frequency domain iteratively as the cache is filled. We illustrate our FreqKV in Figure 2.
@@
For tokens within the context window, the standard attention will be conducted. For tokens out of the window, we will compress the cached KV states in the frequency domain. During each compression step, KV states are converted to the frequency domain via DCT, where low-frequency components are retained before applying IDCT to transform them back to the time domain. Because the retaining size $L$ is smaller than the cache size $N$, subsequent tokens could get in and fill the cache. The compressed cache will be concatenated with KV states of the incoming tokens and compressed together when the cache is filled again. Tokens that appear earlier in the sequence undergo more iterations of compression as the context window expands, whereas less compression will be performed on the more recent tokens. Since the compression is only performed when the sum of cached KV and incoming tokens reaches the preset size, the computation overhead of the compression could be negligible. For example, with the context size of 4096 and the retaining size of 2048, compression is performed every 2048 tokens for contexts exceeding the original window.
@@
Recent work has found the phenomenon of attention sinks that LLMs tend to assign high attention scores to initial tokens (Xiao et al.,, 2024; Han et al.,, 2024). Therefore, we maintain these initial tokens uncompressed in the cache and only compress tokens that come after them.
@@
During training and the prefilling stage of inference, the whole sentence is tokenized and fed into the model. The attention is computed chunk-wise interleaved with the compression operation. After each compression, $N-L-S$ incoming tokens are regarded as a chunk and fill the cache, with $S$ sink tokens uncompressed. As shown in Figure 2, the incoming tokens in each chunk can not attend to the subsequent tokens. The newly incoming token $x_{M}$ will attend to $S$ sink tokens, $L$ compressed “tokens”, $M-N$ previous incoming tokens, and $x_{M}$ itself. The calculation of attention in Equation 7 can be reformulated as follows:
@@
$\mathcal{\widetilde{\displaystyle{\bm{A}}}}(S,N,L,M)=\text{Softmax}\left(\frac{\displaystyle{\bm{q}}_{M}[\displaystyle{\bm{K}}_{0:S-1}\oplus\widetilde{\displaystyle{\bm{K}}}_{0:L-1}^{S:N-1}\oplus\displaystyle{\bm{K}}_{N:M}]^{T}}{\sqrt{d}}\right)\cdot\quad[\displaystyle{\bm{V}}_{0:S-1}\oplus\widetilde{\displaystyle{\bm{V}}}_{0:L-1}^{S:N-1}\oplus\displaystyle{\bm{V}}_{N:M}]$.
@@
It should be noted that key states in LLMs are often equipped with position embeddings like RoPE (Su et al.,, 2023). In FreqKV, key states are compressed and cached before applying RoPE. The cached key states are encoded with RoPE when conducting self-attention. They are allocated with position indices within the cache rather than the original sequence, which results in the context window extension without requiring position extrapolation or interpolation. Compressing key states with RoPE is explored in Appendix B.

### v3

# FreqKV: Key-Value Compression in Frequency Domain for Context Window Extension

@@ Abstract
Existing key-value (KV) cache compression methods for large language models (LLMs) often rely on token eviction, which risks losing critical local information in both long prefilling and decoding scenarios. When extrapolating beyond the pretrained context length, their performance degrades sharply on long-context benchmarks. Motivated by the observation in the frequency domain that the context information is concentrated in the low-frequency components, we propose FreqKV, a parameter-free and architecture-agnostic approach. It iteratively compresses the increasing KV cache in the frequency domain, allowing models to process lengthy contexts efficiently. With minimal training at 8K length, FreqKV extends the context window of LLaMA-2-7B up to 256K tokens while maintaining stable perplexity. Extensive experiments across prefilling and decoding demonstrate that FreqKV enables robust context window extension and consistently outperforms existing KV cache compression methods on LLaMA-2 and LLaMA-3, highlighting its effectiveness for both understanding and generation in long contexts. Our code is available at https://github.com/LUMIA-Group/FreqKV.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable performance in natural language understanding and generation. However, their inference capabilities are fundamentally constrained by a pre-defined context window size established during pre-training. Consequently, LLMs struggle to maintain their performance when processing sequences that exceed the preset context size, significantly limiting their applicability in long-context scenarios.
To alleviate this limitation, several position encoding methods, such as ALiBi (Press et al., 2022), PI (Chen et al., 2023), and LongRoPE (Ding et al., 2024) have been proposed to extend the context window. Relying on full self-attention, the quadratic computational cost renders them prohibitively expensive at scale. LongLoRA (Chen et al., 2024) trains LLMs using shifted sparse attention. Despite training efficiency, their sparse attention fails to be applied during inference, and they still require the original attention on the full sequence.
A popular strategy to improve efficiency is to compress the key-value (KV) cache during inference. Methods such as SnapKV (Li et al., 2024), PyramidKV (Cai. et al., 2024b), and FastKV (Jo et al., 2025) evict tokens deemed less important according to attention scores. However, these compression strategies cannot be extrapolated beyond the context window. The information associated with the evicted tokens will be lost, leading to performance degradation when decoding future tokens.
Another line of work (Zhang et al., 2024; Wan et al., 2025) attempts to merge tokens rather than dropping them, preserving more information. Nevertheless, their performance remains suboptimal without fine-tuning. Concurrently, LoCoCo (Cai et al., 2024) and Activation Beacon (Zhang et al., 2025) introduce additional modules to compress KV states and incorporate the fine-tuned compressing pattern into the decoding procedure of LLMs. However, the increasing memory requirement for additional parameters is inevitable.
In this work, we observe that the KV states of LLMs exhibit strong energy concentration in the low-frequency components of the frequency domain, suggesting that high-frequency parts are largely redundant. Building on this observation, we introduce FreqKV, an efficient context window extension method that iteratively compresses key-value states in the frequency domain. Unlike eviction methods, FreqKV preserves the information of all tokens by transforming KV states into the frequency domain and discarding high-frequency components that carry limited information. By compressing iteratively as the cache grows, FreqKV retains recent tokens with minimal loss while progressively compressing earlier ones. This parameter-free method requires no architectural modifications and can be applied to both the prefilling stage (long-context understanding) and the decoding stage (long-context language modeling and generation).
Extensive experiments show that FreqKV enables robust context window extension. With minimal fine-tuning at 8K length, FreqKV maintains low perplexity as the evaluation context length scales to 256K. It achieves comparable performance to methods that employ full cache or additional compressors in long context language modeling. Experimental results across LongBench, RULER, and Needle-in-a-Haystack indicate that FreqKV surpasses recently studied KV compression methods in the prefilling stage. Furthermore, FreqKV also delivers substantial improvements in long-generation capability on LongGenBench, validating its effectiveness in decoding scenarios. These results establish frequency-domain compression as a viable approach for extending the context window.
@@ 2 Related Work
To alleviate memory and computation costs, a growing body of research explores compressing the KV cache when processing long contexts. One widely studied approach is selective token eviction, which can be traced back to attention sinks (Xiao et al., 2024a; b), suggesting that preserving a small set of initial tokens helps stabilize model performance. Furthermore, MInference (Jiang et al., 2024) propose to accelerate prefilling with unique patterns in long-context attention matrices. More recent methods, SnapKV (Li et al., 2024), InfiniPot (Kim et al., 2024), PyramidKV (Cai. et al., 2024b), GemFilter (Shi et al., 2024), and FastKV (Jo et al., 2025), use attention scores to measure the importance of tokens and discard less important ones. However, LLMs suffer from the permanent loss of the information associated with evicted tokens during decoding. To address this limitation, some researchers introduce cache merging techniques to approximate the original full attention of the existing contexts, such as CaM (Zhang et al., 2024), KVMerger (Wang et al., 2024), and D2O (Wan et al., 2025). Nevertheless, these inference methods often sacrifice performance for efficiency and cannot be extrapolated beyond the context window.
Efficient Context Extension for LLMs.
Recent advancements in context extension for LLMs have focused on efficiently scaling models to handle longer input sequences without significantly increasing computational costs. LongLoRA (Chen et al., 2024) employs shifted sparse attention during the parameter-efficient fine-tuning. However, this sparse attention mechanism is not applicable during inference, necessitating a return to the original full attention post-training. Other techniques, such as LoCoCo (Cai et al., 2024), integrate convolutional operations into LLMs for compressing long contexts. They fine-tune the compression modules together with LLMs. Landmark attention (Mohtashami & Jaggi, 2023) uses landmark tokens to retrieve previous input blocks. Similarly, Activation Beacon (Zhang et al., 2025) introduces a special token to represent the previous context for compression. However, they need a copy of multi-head attention, which amounts to approximately 2B for 7B models. In contrast, our proposed method achieves context extension without introducing any additional parameters.
@@
Learning in the frequency domain is a well-established technique to compress images and accelerate CNNs (Gueguen et al., 2018). It has been observed that CNNs are more sensitive to low-frequency channels than high-frequency channels (Xu et al., 2020). Moreover, Frequency Principle (Xu et al., 2019; Luo et al., 2019) theoretically shows that low frequencies converge faster than high frequencies for deep neural networks. These works have inspired efforts to process natural language. FNet (Lee-Thorp et al., 2022) improves the efficiency of Transformer encoder architectures by replacing the self-attention layers with the Fourier transform to serve the purpose of mixing tokens. Additionally, Fourier Transformer (He et al., 2023) eliminates redundancies in the context through frequency domain processing within encoder architectures.
@@ 3 FreqKV
Figure 1: The average power spectrums of key states and value states in different layers of LLaMA-2-7b. 1000 documents are sampled from CNN/Daily Mail (Hermann et al., 2015). We transform key states and value states to the frequency domain, and average power spectrums over these samples and hidden dimensions.
3.1 Preliminaries
Energy Concentration in the Frequency Domain.
We use DCT (Discrete Cosine Transform) to transform key states and value states from the time domain, which is the sequence dimension, to the frequency domain. The formulations of DCT and the inverse DCT (IDCT) can be referred to in Appendix A.
The average power spectrums in different decoder layers of LLaMA-2-7B are calculated and presented in Figure 1. The figure reveals that the energy of key states and value states is increasingly concentrated in the low-frequency components. The distributions of their power spectrums are similar since they are projected from the same hidden states. While the initial embeddings of natural languages in the first layer exhibit no strong low-frequency bias, subsequent layers gradually shift energy toward low frequencies along the decoding procedure. This observation suggests that the low-frequency components capture most of the essential information, whereas high-frequency parts carry limited signal and can be discarded with minimal loss. Further analysis of frequency components is provided in Appendix B.
Self-Attention with KV Cache.
For the incoming token $x_{N}$, the prefilled $N$ tokens ${\bm{X}}_{0:N-1}=[x_{0},\dots,x_{N-1}]$ are utilized as the cache during decoding. Denote the cached KV states for the previous $N$ tokens ${\bm{X}}_{0:N-1}$ as ${\bm{K}}_{0:N-1}$ and ${\bm{V}}_{0:N-1}$. For simplicity, indices corresponding to layers and heads have been omitted.
Figure 2: Cosine Similarity between perturbed and original inputs for different frequency components of KV states.
3.2 Effect of frequency components
To empirically investigate what information different frequency bands encode, we have conducted a targeted perturbation experiment. We sample 100 CNN/DailyMail documents and introduce controlled perturbations by randomly replacing 1% of words with their synonyms. We then transform the KV states of LLaMA-2 into the frequency domain and compare the cosine similarity between perturbed and original inputs for both low-frequency (lowest 50%) and high-frequency (highest 50%) components. The results across different layers are presented in Figure 2:
It shows that the output states are robust to the perturbation with low-frequency components retained while high-frequency ones are more sensitive, supporting the interpretation that low frequencies encode global information, while high frequencies capture local details.
We further evaluate this distinction on the summarization track of the downstream benchmark LongBench, comparing the performance of llama-2-chat-7b with low and high frequencies retained in Table 1. Summarization requires capturing global semantic structure and long-range coherence. Accordingly, retaining low-frequency components yields substantially better performance.
Table 1: Performance on summarization tasks with high/low-frequency components retained.
Components Retained GovReport QMSum MultiNews
high-frequency 14.21 16.54 15.55
low-frequency 25.51 21.81 26.9
Together, these results strengthen our motivation: low-frequency components primarily encode global semantic contexts and long-range dependencies, whereas high-frequency components encode local details, which explains why preserving low-frequency information is both more robust and more beneficial for long-context understanding.
Figure 3: The overview of our FreqKV. (a) The illustration of the frequency-domain compression. (b) The KV cache will be compressed in an iterative manner to extend the context window. Sink tokens remain uncompressed throughout the process. The tokens after sink tokens will be compressed in the frequency domain and subsequent tokens will continue to get into the cache. When the cache is filled again, the compressed tokens and incoming tokens will be compressed together.
@@ 3.3 KV Compression in the Frequency Domain
To reduce redundancy in the key-value (KV) cache, we compress KV states in the frequency domain as shown in Figure 3a. Specifically, we conduct DCT along the sequence dimension to transfer the KV cache to the frequency domain:
${\bm{Z}}_{0:N-1}^{K}=\text{DCT}(\displaystyle{\bm{K}}_{0:N-1}),\quad\displaystyle{\bm{Z}}_{0:N-1}^{V}=\text{DCT}(\displaystyle{\bm{V}}_{0:N-1}),$ (2)
where ${\bm{Z}}_{0:N-1}^{K}$ and ${\bm{Z}}_{0:N-1}^{V}$ are the spectral representations of the KV cache.
As illustrated in Figure 1, since the low-frequency components exhibit higher magnitudes and carry more information, we will retain them and remove the high-frequency parts for compression. Given a retaining ratio $\gamma$, the cache size is reduced to $L=\gamma\cdot N$ by filtering out NLN-L high-frequency components.
Then we conduct IDCT along the frequency dimension to convert the compressed components back to the time dimension. It should be noted that the signals are normalized by the square root of the component number when applying DCT and IDCT (Appendix A). Therefore, the compressed signals should be rescaled with the factor of $\sqrt{\frac{L}{N}}$ to restore the original amplitude:
$\widetilde{\displaystyle{\bm{K}}}_{0:L-1}^{0:N-1}=\sqrt{\frac{L}{N}}\text{IDCT}(\displaystyle{\bm{Z}}_{0:N-1}^{K}[0:L-1]),\quad\widetilde{\displaystyle{\bm{V}}}_{0:L-1}^{0:N-1}=\sqrt{\frac{L}{N}}\text{IDCT}(\displaystyle{\bm{Z}}_{0:N-1}^{V}[0:L-1]),$ (3)
$\widetilde{\displaystyle{\bm{K}}}_{0:L-1}^{0:N-1}$ and $\widetilde{\displaystyle{\bm{V}}}_{0:L-1}^{0:N-1}$ are the compressed KV cache of size $L$ in the time domain. The superscript “0:N1N-1” indicates the original token range, while the subscript “0:L1L-1” means the retaining size. The incoming token $x_{N}$ will attend to the compressed KV cache as follows:
@@ 3.4 Context Extension via Iterative Compression
Extending the context window of LLMs is fundamentally constrained by memory and computation cost. To address this, FreqKV employs an iterative compression strategy in the frequency domain that constrains the effective cache size while enabling processing of arbitrarily long sequences. The overall pipeline is illustrated in Figure 3b.
@@
For tokens within the context window, the standard attention will be conducted. Once the cache reaches the maximum size, we will compress it in the frequency domain. Specifically, cached KV states are converted to the frequency domain via DCT, where low-frequency components are retained before applying IDCT to transform them back to the time domain. The incoming tokens are appended to the compressed cache, and the process repeats when the cache is filled again. As a result, earlier tokens experience more iterations of compression as the context window expands, whereas less compression will be performed on the more recent tokens—a property well aligned with the autoregressive nature of LLMs. The compression is triggered only once every NLN-L tokens, with the complexity of $O(NlogN)$ for each compression. Therefore, the computational overhead of the compression could be negligible.
@@
Recent work has found the phenomenon of attention sinks that LLMs tend to assign high attention scores to initial tokens (Xiao et al., 2024b; Han et al., 2024). Therefore, we maintain these initial tokens uncompressed in the cache and only compress tokens that come after them.
@@
During training and prefilling, FreqKV will process the sentence chunk-wise, interleaving attention computation with the compression operation. After each compression, NLSN-L-S incoming tokens form a chunk that fills the cache, while $S$ sink tokens remain uncompressed, and $L$ compressed “tokens” represent earlier history. Consequently, the newly incoming token $x_{M}$ will attend to $S$ sink tokens, $L$ compressed tokens, MNM-N previous incoming tokens, and itself:
@@
$\mathcal{\widetilde{\displaystyle{\bm{A}}}}(S,N,L,M)=\text{Softmax}\left(\frac{\displaystyle{\bm{q}}_{M}[\displaystyle{\bm{K}}_{0:S-1}\oplus\widetilde{\displaystyle{\bm{K}}}_{0:L-1}^{S:N-1}\oplus\displaystyle{\bm{K}}_{N:M}]^{T}}{\sqrt{d}}\right)\cdot[\displaystyle{\bm{V}}_{0:S-1}\oplus\widetilde{\displaystyle{\bm{V}}}_{0:L-1}^{S:N-1}\oplus\displaystyle{\bm{V}}_{N:M}]$,
where $\widetilde{\displaystyle{\bm{K}}}_{0:L-1}^{S:N-1}$and $\widetilde{\displaystyle{\bm{V}}}_{0:L-1}^{S:N-1}$ denote the compressed KV cache of size $L$ for tokens ${\bm{X}}_{S:N-1}=[x_{S},\dots,x_{N-1}]$.
@@
It should be noted that key states in LLMs are often equipped with position embeddings like RoPE (Su et al., 2024). In FreqKV, key states are compressed and cached before applying RoPE. The cached key states are then encoded with RoPE when conducting self-attention. They are allocated with position indices within the cache rather than the original sequence, which results in the context window extension without requiring position extrapolation or interpolation. Further discussion of RoPE integration is provided in Appendix C.